OpenStudy (anonymous):

Which set represents the domain of the function shown? {(−3, 6), (0, 2), (4, 7), (11, 15)} {(6, −3), (2, 0), (7, 4), (15, 11)} {2, 6, 7, 15} {−3, 0, 4, 11} {−3, 0, 2, 4, 6, 7, 11, 15} Given that h(x) = 3x −19, find the value of x that makes h(x) = 71.

OpenStudy (anonymous):

they are two separate questions. please help

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hey :)

OpenStudy (misty1212):

for the domain of \[\{(−3, 6), (0, 2), (4, 7), (11, 15)\} \]list the first numbers in each pair

OpenStudy (anonymous):

i always mix it up with the range lol. now i understand that one thank you!

OpenStudy (misty1212):

for the second one, solve \[3x-19=71\]

OpenStudy (anonymous):

and then that will be the answer?

OpenStudy (misty1212):

add 19 and then divide by 3 what do you get ?

OpenStudy (anonymous):

30

OpenStudy (misty1212):

yup

OpenStudy (anonymous):

thank you! could you help me with a few more? @misty1212

OpenStudy (misty1212):

sure

OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (2, 4). y = 2 x = 2 y = 4 x = 4 Write the equation of the line that passes through the points (3, 6) and (5, 18) using function notation. f(x) = 2x + 12 f(x) = 6x − 12 y = 2x + 12 y = 6x − 12 Choose the equation below that represents the line that passes through the point (7, −2) and has a slope of −3. y − 2 = −3(x + 7) y + 7 = −3(x − 2) y + 2 = −3(x − 7) y − 7 = −3(x + 2)

OpenStudy (misty1212):

Choose the equation of the horizontal line that passes through the point (2, 4). horizontal line looks like \(y=\text{something}\) in this case the "something" is 4

OpenStudy (anonymous):

y = 4?

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

yay!! ok :)

OpenStudy (misty1212):

Write the equation of the line that passes through the points (3, 6) and (5, 18) using function notation. slope is \(\frac{18-6}{5-3}=\frac{12}{2}=6\)

OpenStudy (misty1212):

\[ f(x) = 6x − 12\]is the only choice written in function notation

OpenStudy (anonymous):

oh ok that makes sense :)

OpenStudy (misty1212):

Choose the equation below that represents the line that passes through the point (7, −2) and has a slope of −3. point slope formula \[y-y_1=m(x-x_1)\] in this case \[y+2=-3(x-7)\]

OpenStudy (anonymous):

thank you sooooo much

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]